OpenStudy (anonymous):

A 36 inches piece of string is cut into two pieces. one piece is used to form a circle while the other is used to form a square, how should the string be cut so the sum of the areas is a minimum?

OpenStudy (anonymous):

You still there?

OpenStudy (kesumonu):

u can take x is the piece taken out of 36 inch string now u have two string of length x, (36-x) let 1st one is used to make square and 2nd for circle so 4a=x,=> a=x/4; Asquare=x^2/16; for circle 2(pi)r=(36-x) Acircle=(36-x)^2/4pi now A=(x^2/16)+((36-x)^2/4pi) take dA/dx=0 u will get x=144/(4+pi)=20.16 (aprox) check d^2A/dx^2 whether +ve or -ve here it comes +ve so our x=20.16 is a point of minima and hence this will give us minimum area